OpenStudy (anonymous):

In stormy contidions, a delivery turck can travel a route in 14 hr. In good weather, the same trip ccan be made in 10.4 hr because the truck tralvels 15km/hr faster. Find the speed of the truck in stormy weather and the speed in good weather

OpenStudy (anonymous):

has less wind resistance

OpenStudy (anonymous):

do you know where to start?

OpenStudy (anonymous):

@Jlynn89 I'm trying to figure out where you need help how far have you gotten?

OpenStudy (anonymous):

I don't even know how to start it.

OpenStudy (anonymous):

okay no worries. The key to problems like this is to let one of the things you are trying to find be "x"

OpenStudy (anonymous):

so we will let the slower speed be x for us

OpenStudy (anonymous):

can you write the faster speed in terms of x?

OpenStudy (anonymous):

I do not know what you mean by that. The "terms of x" bit

OpenStudy (anonymous):

this is not the correct anwer for writing the faster speed in terms of x but it should explain what I mean by in terms of x: faster = x +5

OpenStudy (anonymous):

let me put it this way how much faster is the good weather speed compared to the bad weather speed?

OpenStudy (anonymous):

Oh okay got you.

OpenStudy (anonymous):

so what did you get?

OpenStudy (anonymous):

What you meant by terms of x, I didn't see you had writen something else. Sorry I am really crap at this section.

OpenStudy (anonymous):

well I mean what did you get when you wrote the faster speed in terms of x? I did not give you the correct answer above

OpenStudy (anonymous):

" truck tralvels 15km/hr faster"

OpenStudy (anonymous):

Ok I set it up like what is in the picture then did 14x= 10.5 (x+75) then distributed 14x= 10.5x+157.5 subtracted 10.5 from itself and tehn from 14 3.5x=157.5 divide 3.5 on both sides and ended up with x= 45 x=45 x+15 ended up being 45+15 and got 60

OpenStudy (anonymous):

seems right to me :)

OpenStudy (anonymous):

lol thank you! I get really confused with this ;_;

OpenStudy (anonymous):

not a problem just always remember to see if you can rewrite one of the numbers you are looking for in terms of the other number you are looking for